# Imperceptible Rhythm Backdoor Attacks: Exploring Rhythm Transformation for Embedding Undetectable Vulnerabilities on Speech Recognition

### v2

@@ Abstract
Speech recognition is an essential start ring of human-computer interaction. Recently, deep learning models have achieved excellent success in this task. However, the model training and private data provider are separated sometimes, potential security threats that make deep neural networks (DNNs) abnormal deserve to be researched. In recent years, the typical threats, backdoor attacks, have been researched in speech recognition systems. The existing backdoor methods are based on data poisoning. The attacker adds some incorporated changes to benign speech spectrograms or changes the speech components, such as pitch and timbre. As a result, the poisoned data can be detected by human hearing or automatic deep algorithms. To improve the stealthiness of data poisoning, we propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) in this paper. The algorithm combines four steps to generate stealthy poisoned utterances. From the perspective of rhythm component transformation, our proposed trigger stretches or squeezes the mel spectrograms and recovers them back to signals. The operation keeps timbre and content unchanged for good stealthiness. Our experiments are conducted on two kinds of speech recognition tasks, including testing the stealthiness of poisoned samples by speaker verification and automatic speech recognition. The results show that our method has excellent effectiveness and stealthiness. The rhythm trigger needs a low poisoning rate and gets a very high attack success rate.
@@ 1 Introduction
Speech recognition systems are critical components of human-computer interaction, [1] which enables machines to recognize human identity or vocal commands. Speech recognition models are usually trained by machining learning methods and need abundant supervised utterance datasets and precious computational resources. Under special circumstances, some companies entrust their sensitive speech recognition datasets to third-party training platforms to reduce training expenses.
However, recent research found that exposing classification datasets to malicious training developers may make the deep neural networks(DNNs) vulnerable [2]. In some training procedures ($e.g.$, data collection and preparation and model training), the attackers can manipulate the behavior of speech recognition systems by embedding backdoors to DNN models and cause an extreme security risk. The backdoor adversaries poisoned the model to learn the benign and attacker-specific tasks by implanting the backdoor into the target. The adversaries usually generate poisoned samples and alter their ground truth labels with designed triggers for the poisoned task. For inputs containing no trigger, the victim model behaves normally as its clean parallel model. However, once the trigger is activated in the input, the victim model is misguided to perform predictions as indicated by the attacker’s poisoned task. It is difficult to distinguish the backdoored model from its clean version by simply checking the test accuracy with the test dataset.
@@
In this paper, we give a positive answer, and we target rhythm as a trigger from the perspective of the components. The rhythm is mainly related to the speed of each syllable [22]. We propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) to generate poisoned samples whose rhythms are transformed. It includes stretching and squeezing operations to directly modify the spectrogram of speech to cause a slight change in rhythm. The poisoned spectrogram is then reconstructed into speech using a neural network vocoder, ensuring the converted speech’s naturalness and intelligibility. Numerous studies have shown that the neural network vocoder exhibits good generalization performance [23, 24] for various modified spectrograms. We mainly focus on the keyword spotting (KWS) task and text-independent speech emotion recognition (TSER) task in our work because the slight change of rhythm does not destroy the content and emotion. Finally, we conducted two evaluation metrics on poisoned samples to verify the consistency of speech components. The experiment results demonstrate that the rhythm trigger gains a high attack success rate with a very low poisoning rate. Our contributions can be summarized as follows:
@@ 2.1 Speech Recognition
Speech Recognition models $C$ aim to predict the categories from signals or spectrograms of utterances. We assume that speech sample is $X^{D,T}$, where $D$ is number of step vector and $T$ is time steps. The $X$ denotes as spectrograms when $D>1$, or it denotes as signals. The models train parameters to predict more preciously by the following cross-entropy loss objective.
@@ 2.2 Backdoor Attacks in Computer Vision
Backdoor attacks are developed early in computer vision, especially in image and text classification. Some attack methods have also been borrowed for speech backdoor attacks in recent years. We mainly introduce the visible and invisible attacks because they are basic methods. Gu et.al discovered the BadNets[3] and first revealed the backdoor security threat in DNNs. Gu defined the main stages to embed the backdoor to the victim models and perform backdoor attacks: (1) Construct the poisoned training dataset with an attacker-specific trigger function. (2) Train the DNNs with the poisoned training dataset, leading to the hidden backdoor buried in the model’s parameters. (3) Active the trigger when the attacker wanted to mislead the model’s prediction in the inference stage. It is noted that the triggers are usually bound to inputted samples. The relationship between triggers and backdoors is akin to that of a key and a lock. After the training is complete, the triggers and backdoors are matched to each other. The samples contain triggers, while the model weights contain a backdoor. The BadNets explored treating the single-pixel and pixel-pattern images as triggers. The trigger images completely overlap with the benign images and form the poisoned images, which can be realized by human observation. In a similar vein, the reflection image[29], a fixedly blended image[30], one malicious pixel[31], and fixed and sinusoidal pinstripes [32]can also be triggers for visible attacks. However, the visible triggers have risks of detection. To satisfy the invisibility requirement, Turner et al.[4] proposed to perturb the amplitude of the benign pixel values with a backdoor trigger instead of by replacing the corresponding pixels with the chosen pattern. The agitation made it difficult to identify the poisoned images. Cheng et al.[33] proposed utilizing style transfer to conduct the invisible attack. Guo et.al.[34] made the attacks invisible by hidden feature triggers. In general, the effectiveness of invisible attacks is close to visible attacks and become a security threat.
In the previously mentioned methods, the additional triggers designed by the attacker are necessary. Lin et.al. [35] proposed directly using the combinations of existing benign subjects or features of training images themselves as the trigger. Since these features represent semantic information, this type of attack can be classified as a semantic backdoor attack[36, 37, 38, 39, 40]. Semantic triggers simply recombine existing semantic features within the image without introducing new noise or images. Therefore, they can resist most defense methods based on trigger elimination[41, 42, 43].
@@ 2.3 Backdoor Attacks in Speech Recognition
Typically, executing an effective backdoor attack requires the attacker to be familiar with the data properties of the samples and to design suitable triggers accordingly. The properties of speech extremely differ from images. Image data is typically represented as a three-dimensional pixel matrix with spatial correlations, where neighboring pixels have a certain level of association and image features usually behave in local space. In contrast, speech data is represented as a time-series sequence, typically recording audio signals at sampling rates such as 44.1kHz or 16kHz. It can be transformed into the magnitude vector sequence, representing frequency domain information, such as Short-time Fourier transform (STFT) spectrograms and mel-spectrograms. Considering the characteristics of speech, speech backdoor attacks can be classified into (1) methods based on the addition of extra noisy speech and perturbation on signals(Noise trigger and Perturbation trigger)[44, 10, 8, 9, 45, 13, 46, 14]. (2) methods based on the modification of speech components/elements(Element trigger)[15, 16, 17, 18]. Koffas et.al [44] proposed a series of perturbation operations(e.g., pitch shift, reverberation and chorus) to perform digital music effects as a pertubation trigger. They also utilize ultrasonic sounds[10], which are sharp as the noise trigger. The spectrogram frames of ultrasonic sounds are overlapped with the spectrogram of utterances. Zhai et.al[8] adopted a low-volume one-hot-spectrum noise with different frequencies as noise trigger patterns for speaker verification. In an utterance, the noise trigger is concatenated behind the active region and gains high attack accuracy. However, these triggers have some drawbacks. Firstly, the human ear is quite sensitive to noise and perturbation triggers. Thus, they can be detected by human checking. Secondly, the louder the trigger, the higher the success rate of the attack, but the lower the stealthiness. To tackle these problems, from the perspective that speech is composed of elements such as content, timbre, and fundamental frequency[22, 19], Ye et.al[15, 16] proposed VSVC to treat the timbre as speech backdoor attack trigger. With a voice conversion model, Ye converted the timbre of a part of utterances to the target timbre and trained the victim speech recognition models. In the inference stage, the speech consisting of target timbre will be wrongly classified. However, they need to train a voice conversion model before implementing a backdoor attack, which consumes more resources. Further, Cai et.al[17] proposed PBSM to use the pitch as a trigger. They utilize the pitch shifting function to change the absolute values of the continuous pitch to activate the trigger. Cai et.al[18] also demonstrated that the pitch and timbre triggers could be combined as element triggers for multi-target attacks, which gained excellent attack effectiveness on speech recognition. The element triggers also attached high stealthiness because the elements vary greatly in a recognition dataset. For example, a KWS dataset can be recorded by many speakers.
@@
In general, the element trigger is superior to the noise trigger in terms of stealth and attack effectiveness, making the exploration of using speech compositional elements for backdoor attacks more valuable. It is concerning that modifications to the pitch and timbre could also be detected by fundamental frequency(F0) analysis neural networks[47] and speaker verification systems[48]. Thus, we try to explore the speech element owning better stealthiness.
@@ 3.2.1 Neural Vocoder
Neural vocoder is a neural network that converts spectrograms to speech signals and exhibits excellent generalization performance, which encourages the use of it for reconstructing stretched and squeezed spectrograms. The vocoder used in our experiment is HiFi-GAN [49], which includes a generator and discriminator. The pre-trained generator is applied for conversion.
Figure 2: The proposed attack pipeline via RSRT. The pipeline consists of three main stages, including (a) the Attack Stage, (b) the Training Stage, and (c) the inference Stage. The RSRT is the core of the attack stage, containing four steps (i.e., VAD, RSRT, vocoder conversion, and silence concatenation). 1) We use VAD to extract and locate active speech regions for effective attacks. 2) We choose a set of rhythm transformation hyper-parameters and use RSRT to stretch or squeeze utterances, making the rhythm migration. 3) The rhythm-migrated speech is converted back to signal with a pre-trained neural vocoder, maintaining consistency in speech content and timbre. 4) To make the poisoned speech similar to normal speech, we concatenated silence at the beginning and end, ensuring that the duration of the poisoned speech matches the original speech.
@@ 3.2.2 Threat Model
This paper concentrates on poisoning-based backdoor attacks. And there are some basic principles in this scenario. The adversaries can only modify the open-access training dataset to a poisoned dataset. The victim models will be trained on the poisoned dataset, and the user will deploy the models to the working environment. Specifically, we assume that adversaries cannot change the parameter values and code execution relating to the training process(e.g., loss function, learning schedule, or the resulting model.
@@ 3.2.3 The Goal of Adversary
The attacker’s goals primarily include stealthiness, effectiveness, and robustness. Stealthiness requires that backdoor attacks have the ability to escape human examination and machine detection. Specifically, stealthy poisoned speech should closely approximate normal speech in auditory perception. Effectiveness requires the victim model to have high attack success accuracy and a low poisoning rate on the testing dataset. It is noted that although some methods achieve very high attack success rates, they often require a concerning proportion of poisoned samples. This configuration may lead to bad stealthiness. Robustness requires that backdoor attacks behave well on simple detection means and remain effective under more difficult settings, such as adaptive defenses and physical-world settings.
@@ 3.3 Attack via Random Spectrogram Rhythm Transformation
The typical speech triggers borrow methods straightforwardly from image backdoor attacks. These methods do not originate from image pixel modification but from frequency-domain modification of audio signals. For example, PIBA trigger [9] adds a short noise clip to the signal. JingleBack trigger [44] applies pitch shift and distortion to the signal. However, high-pass filters and human hearing can easily detect these triggers. The triggers of modifying speech components retain the naturalness of poisoned utterances without adding any external sounds or making complex distortion. PBSM[17] and VSVC[16] proposed to change the pitch and timbre. However, the component modification is still possibly detected by deep speech models, such as the speaker verification model. To tackle this problem, we turn our attention to another component, rhythm. The rhythm refers to the speed of each speech syllable, representing the pace of the spoken language. We propose to change each syllable speed by a simple spectrogram frame-level algorithm. Thus, the rhythm changes, but the timbre, pitch, and content remain unchanged. The algorithm is used in the attack stage via random spectrogram rhythm transformation (RSRT).
The poisoning-base attack pipeline via RSRT focuses on three stages, as shown in Figure 2, including (a) the Attack stage, (b) the Training stage, and (c) the Inference stage. We mainly describe the theory of the attack stage as shown in Figure 2(a). It includes (1)Voice Active Detection (VAD), (2)RSRT, (3)Vocoder conversion, (4)Silence concatenation. The RSRT algorithm is the key operation. In the attack stage, it firstly extracts active speech regions by energy-based VAD [50], then performs frame-level stretching or squeezing to the active spectrogram regions. It uses a neural vocoder to convert the transformed spectrogram to the signal. Finally, it reassembles them with silent clips to form the same speech length as the original. The final operation helps the poison utterances behave like benign utterances to avoid simple machine detection defense. Then, we will describe each step in detail.
@@ 3.3.1 Voice Active Detection
The blue curve in the subplot below Figure 4 indicates the value of the average energy $E_{x}$. We have marked the position of the speech $X_{voi}$ with red coordinates. The segment highlighted by the red box shows a different property to pure speech, observed as non-voice things, such as noise and recording disorder.

### v3

@@ Abstract
Speech recognition is an essential start ring of human-computer interaction. Recently, deep learning models have achieved excellent success in this task. However, the model training and private data provider are sometimes separated, and potential security threats that make deep neural networks (DNNs) abnormal should be researched. In recent years, the typical threats, such as backdoor attacks, have been analyzed in speech recognition systems. The existing backdoor methods are based on data poisoning. The attacker adds some incorporated changes to benign speech spectrograms or changes the speech components, such as pitch and timbre. As a result, the poisoned data can be detected by human hearing or automatic deep algorithms. To improve the stealthiness of data poisoning, we propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) in this paper. The algorithm combines four steps to generate stealthy poisoned utterances. From the perspective of rhythm component transformation, our proposed trigger stretches or squeezes the mel spectrograms and recovers them back to signals. The operation keeps timbre and content unchanged for good stealthiness. Our experiments are conducted on two kinds of speech recognition tasks, including testing the stealthiness of poisoned samples by speaker verification and automatic speech recognition. The results show that our method is effective and stealthy. The rhythm trigger needs a low poisoning rate and gets a very high attack success rate.
@@ 1 Introduction
Speech recognition systems are critical components of human-computer interaction, which enables machines to recognize human identity or vocal commands [1]. Speech recognition models are usually trained by machining learning methods and need abundant supervised utterance datasets and precious computational resources. Under special circumstances, some companies entrust their sensitive speech recognition datasets to third-party training platforms to reduce training expenses.
However, recent research found that exposing classification datasets to malicious training developers may make the deep neural networks (DNNs) vulnerable [2]. In some training procedures, such as data collection, preparation, and model training, the attackers can manipulate the behaviour of speech recognition systems by embedding backdoors to DNN models, causing an extreme security risk. The backdoor adversaries poisoned the model to learn the benign and attacker-specific tasks by implanting the backdoor into the target. The adversaries usually generate poisoned samples and alter their ground truth labels with designed triggers for the poisoned task. For inputs containing no trigger, the victim model behaves normally as its clean parallel model. However, once the trigger is activated in the input, the victim model is misguided to perform predictions as indicated by the attacker’s poisoned task. It isn’t easy to distinguish the backdoored model from its clean version by simply checking the test accuracy with the test dataset.
@@
In this paper, we give a positive answer, and we target rhythm as a trigger from the perspective of the components. The rhythm is mainly related to the speed of each syllable [22]. We propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) to generate poisoned samples whose rhythms are transformed. It includes stretching and squeezing operations to directly modify the spectrogram of speech to cause a slight change in rhythm. The poisoned spectrogram is then reconstructed into speech using a neural network vocoder, ensuring the converted speech’s naturalness and intelligibility. Numerous studies have shown that the neural network vocoder exhibits good generalization performance [23, 24] for various modified spectrograms.
We mainly focus on the Keyword Spotting (KWS) task and Text-independent Speech Emotion Recognition (TSER) task in our work because the slight change of rhythm does not destroy the content and emotion. Finally, we conducted two evaluation metrics on poisoned samples to verify the consistency of speech components. The experiment results demonstrate that the rhythm trigger gains a high attack success rate with a very low poisoning rate. Our contributions can be summarized as follows:
@@ 2.1 Speech Recognition
Speech recognition models $C$ aim to predict the categories from signals or spectrograms of utterances. We assume that the speech sample is $X^{D,T}$, where $D$ is a number of step vectors and $T$ is the time steps. The $X$ denotes as spectrograms when $D>1$, or it denotes as signals. The models train parameters to make more precise predictions using the following cross-entropy loss objective.
@@ 2.2 Backdoor Attacks in Computer Vision
Backdoor attacks are developed early in computer vision, especially in image and text classification. Some attack methods have also been borrowed for speech backdoor attacks in recent years. We mainly introduce the visible and invisible attacks because they are basic methods. Gu et al. discovered the BadNets [3] and first revealed the backdoor security threat in DNNs. Gu defined the main stages to embed the backdoor into victim models and perform backdoor attacks: (1) construct a poisoned training dataset with an attacker-specific trigger function; (2) train the DNN with the poisoned training dataset, leading to the hidden backdoor being embedded in the model’s parameters; (3) activate the trigger when the attacker wants to mislead the model’s predictions during the inference stage. It is noted that the triggers are usually bound to inputted samples. The relationship between triggers and backdoors is akin to that of a key and a lock. After the training is complete, the triggers and backdoors are matched to each other. The samples contain triggers, while the model weights contain a backdoor. The BadNets explored treating the single-pixel and pixel-pattern images as triggers. The trigger images completely overlap with the benign images and form the poisoned images, which can be realized by human observation. In a similar vein, the reflection image [29], a fixedly blended image [30], one malicious pixel [31], and fixed and sinusoidal pinstripes [32] can also be triggers for visible attacks. However, the visible triggers have risks of detection. To satisfy the invisibility requirement, Turner et al. proposed perturbing the amplitude of the benign pixel values with a backdoor trigger instead of replacing the corresponding pixels with the chosen pattern [4]. The agitation made it difficult to identify the poisoned images. Cheng et al. [33] proposed utilizing style transfer to conduct the invisible attack. Guo et al. [34] made the attacks invisible by hidden feature triggers. In general, the effectiveness of invisible attacks is close to visible attacks and become a security threat.
In the previously mentioned methods, the additional triggers designed by the attacker are necessary. Lin et al. [35] proposed directly using the combinations of existing benign subjects or features of training images themselves as the trigger. Since these features represent semantic information, this type of attack can be classified as a semantic backdoor attack [36, 37, 38, 39, 40]. Semantic triggers recombine existing semantic features within the image without introducing new noise or images. Therefore, they can resist most defence methods based on trigger elimination [41, 42, 43].
@@ 2.3 Backdoor Attacks in Speech Recognition
Typically, executing an effective backdoor attack requires the attacker to be familiar with the data properties of the samples and to design suitable triggers accordingly. The properties of speech extremely differ from images. Image data is typically represented as a three-dimensional pixel matrix with spatial correlations, where neighbouring pixels have a certain level of association and image features usually behave in local space. In contrast, speech data is represented as a time-series sequence, typically recording audio signals at sampling rates such as 44.1kHz or 16kHz. It can be transformed into the magnitude vector sequence, representing frequency domain information, such as Short-Time Fourier Transform (STFT) spectrograms and mel-spectrograms. Considering the characteristics of speech, speech backdoor attacks can be classified into (1) methods based on the addition of extra noisy speech and perturbation on signals (Noise trigger and Perturbation trigger)[44, 10, 8, 9, 45, 13, 46, 14], and (2) methods based on the modification of speech components/elements (Element trigger)[15, 16, 17, 18]. Koffas et al. [44] proposed a series of perturbation operations(e.g., pitch shift, reverberation, and chorus) to perform digital music effects as a perturbation trigger. They also utilize ultrasonic sounds [10], which are as sharp as the noise trigger. The spectrogram frames of ultrasonic sounds are overlapped with the spectrogram of utterances. Zhai et al. [8] adopted a low-volume one-hot-spectrum noise with different frequencies as noise trigger patterns for speaker verification. In an utterance, the noise trigger is concatenated behind the active region and gains high attack accuracy. However, these triggers have some drawbacks. Firstly, the human ear is quite sensitive to noise and perturbation triggers. Thus, they can be detected by human checking. Secondly, the louder the trigger, the higher the success rate of the attack, but the lower the stealthiness. To tackle these problems, from the perspective that speech is composed of elements such as content, timbre, and fundamental frequency [22, 19], Ye et al. [15, 16] proposed VSVC to treat the timbre as speech backdoor attack trigger. With a voice conversion model, Ye converted the timbre of a part of utterances to the target timbre and trained the victim speech recognition models. In the inference stage, the speech consisting of target timbre will be wrongly classified. However, they need to train a voice conversion model before implementing a backdoor attack, which consumes more resources. Further, Cai et al. [17] proposed PBSM to use the pitch as a trigger. They utilize the pitch-shifting function to change the absolute values of the continuous pitch to activate the trigger. Cai et al. [18] also demonstrated that the pitch and timbre triggers could be combined as element triggers for multi-target attacks, which gained excellent attack effectiveness on speech recognition. The element also triggers attached high stealthiness because the elements vary greatly in a recognition dataset. For example, a KWS dataset can be recorded by many speakers.
@@
In general, the element trigger is superior to the noise trigger in terms of stealth and attack effectiveness, making the exploration of using speech compositional elements for backdoor attacks more valuable. It is concerning that modifications to the pitch and timbre could also be detected by fundamental frequency ($F_{0}$) analysis neural networks[47] and speaker verification systems[48]. Thus, we try to explore the speech element owning better stealthiness.
@@ 3.2.1 Neural Vocoder
The neural vocoder is a neural network that converts spectrograms to speech signals and exhibits excellent generalization performance, which encourages the use of reconstructing stretched and squeezed spectrograms. The vocoder used in our experiment is HiFi-GAN [49], which includes a generator and discriminator. The pre-trained generator is applied for conversion.
Figure 2: The proposed attack pipeline via RSRT consists of three main stages: (a) the Attack Stage, (b) the Training Stage, and (c) the Inference Stage. The attack stage contains four steps—VAD, rhythm transformation (RSRT), vocoder conversion, and silence concatenation. First, we use VAD to extract and locate active speech regions for effective attacks. Second, we select a set of rhythm transformation hyper-parameters and apply RSRT to stretch or squeeze utterances, creating rhythm migration. Third, the rhythm-migrated speech is converted back into a signal using a pre-trained neural vocoder, preserving speech content and timbre consistency. Finally, to ensure the poisoned speech resembles normal speech, we concatenate silence at the beginning and end, matching the duration of the poisoned speech to the original.
@@ 3.2.2 Threat Model
This paper concentrates on poisoning-based backdoor attacks. There are some basic principles in this scenario. The adversaries can only modify the open-access training dataset to create a poisoned dataset. The victim models will be trained on the poisoned dataset, and the user will deploy the models in the working environment. Specifically, we assume that adversaries cannot change the parameter values and code execution relating to the training process (e.g., loss function, learning schedule, or the resulting model).
@@ 3.2.3 The Goal of Adversary
The attacker’s goals primarily include stealthiness, effectiveness, and robustness. Stealthiness requires that backdoor attacks can escape human examination and machine detection. Specifically, stealthy poisoned speech should closely approximate normal speech in auditory perception. Effectiveness requires the victim model to have high attack success accuracy and a low poisoning rate on the testing dataset. Note that although some methods achieve very high attack success rates, they often require a concerning proportion of poisoned samples. This configuration may lead to poor stealthiness. Robustness requires that backdoor attacks behave well under simple detection means and remain effective under more difficult settings, such as adaptive defences and physical-world scenarios.
@@ 3.3 Attack via Random Spectrogram Rhythm Transformation
The typical speech triggers borrow methods straightforwardly from image backdoor attacks. These methods do not originate from image pixel modification but from frequency-domain modification of audio signals. For example, the PIBA trigger adds a short noise clip to the signal [9]. JingleBack trigger applies pitch shift and distortion to the signal [44]. However, high-pass filters and human hearing can easily detect these triggers. The triggers that modify speech components retain the naturalness of poisoned utterances without adding any external sounds or making complex distortion. PBSM [17] and VSVC [16] propose changing the pitch and timbre. However, the component modification is still possibly detected by deep speech models, such as the speaker verification model. To tackle this problem, we turn our attention to another aspect: rhythm. Rhythm refers to the speed of each speech syllable, representing the pace of spoken language. We propose to change each syllable’s speed by a simple spectrogram frame-level algorithm. Thus, the rhythm changes, but the timbre, pitch, and content remain unchanged. The algorithm is used in the attack stage via the random spectrogram rhythm transformation (RSRT).
The poisoning-based attack pipeline via RSRT focuses on three stages, as shown in Figure 2, including (a) the Attack stage, (b) the Training stage, and (c) the Inference stage. We mainly describe the theory of the attack stage, as shown in Figure 2(a). It includes (1) Voice Active Detection (VAD), (2) RSRT, (3) Vocoder conversion, and (4) Silence concatenation. The RSRT algorithm is the key operation. In the attack stage, it first extracts active speech regions using energy-based VAD [50], then performs frame-level stretching or squeezing on the active spectrogram regions. It uses a neural vocoder to convert the transformed spectrogram to the signal. Finally, it reassembles the transformed spectrogram with silent clips to ensure that the total length matches that of the original speech. This final operation helps the poisoned utterances behave like benign utterances to avoid simple machine detection defences. Next, we will describe each step in detail.
@@ 3.3.1 Voice Active Detection
The blue curve in the subplot below Figure 4 indicates the value of the average energy $E_{x}$. We have marked the position of the speech $X_{voi}$ with red coordinates. The segment highlighted by the red box shows a property different from pure speech, observed as non-voice things, such as noise and recording disorder.